Calculate the decay constant (lambda) of radioactive materials of half-life 10 ns?

Will rate thank you!
Answer #1

N= N_0 e^ { -\lambda t}

At half-life, t= T_{\frac{1}{2}}, \ N= \frac{N_0}{2}

\frac{N_0}{2} = N_0 e^{-\lambda T_{\frac{1}{2}}} \Rightarrow \lambda T_{\frac{1}{2}}=ln(2)\newline \lambda= \frac{ln(2)}{T_{\frac{1}{2}}} = \frac{ln(2)}{10 \times 10^{-9}} =0.0693 \times 10^{9} / s = 6.93 \times 10^{7} / s \ (Ans)